a constant volume chamber of 0.3 m3 capacity contains 2kg of gas at 15 degree celcius. Heat is transferred to the gas until the temperature is 100 degree celcius. Find the work done, the heat transferred, the changes in internal energy and enthalpy. Take Cp = 1.97 Kj/Kg K, and Cv = 1.51 Kj/Kg K.

Grade:12th pass

1 Answers

Yuvraj Singh
2 years ago
Characteristics gas constant R is; R =Cp-Cy R 1.968-1.507 R = 0.461 KJ/kgK Universal gas constant Ry is; Ry = 8.314 KJ/kg - mole K Ry = MR So, Molecular weight M is; M 8.51/o461 M = 18.03 kg- mole T = 5°C 5+273 278 K T2 100°C = 100 +273 = 373 K In Constant volume process i.e. Isochoric process Work done is zero. So; Workdone dW = 0
2. Heat transfer dQ = m Cy (T2-Ti) = 2 X 1.507 X (373 - 278) = 286.33 K 3. Change in Internal energy